A rectangle has an area of 180 centimeters squared and a perimeter of 54 centimeters. What are its dminesions?
Answer by mananth(16946)

let length be L & width be W
2(L+W) = 54
L+W = 27
LW = 180
L= 180/W
..
180/W + W = 27
(180+W^2)/W = 27
180+W^2 = 27W
W^2-27W+180=0
W^2-15W-12W+180=0
W(W-15)-12(W-15)=0
(W-15)(W-12)=0
W= 15 OR 12
The dimensions are 15cm by 12 cm
..
m.ananth@hotmail.ca